# Optically Driven Orbital Hall Transport in Floquet Odd-Parity Collinear Altermagnets with High Chern Numbers

### 12 Mar 2026 v1

@@ Abstract
Introduction-Altermagnets provide a unique and significant platform for spin-polarized phenomena [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], distinguished by their nonrelativistic spin splitting in k-space protected by crystalline symmetries [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Typically, Altermagnets exhibit even-parity spin splitting, giving rise to Fermi surfaces with d-, g-, and i-wave symmetries [26, 27]. In contrast, magnets hosting unconventional odd-parity spin splitting (e.g., p- and f-wave types) are of particular significance due to their intimate connection with unconventional superfluidity [28, 29, 30]. However, odd-parity spin splitting has been largely limited to noncollinear magnets [31, 32, 33, 34, 35], whose complex magnetic textures are often fragile against external perturbations, restricting their practical applicability. Realizing odd-parity altermagnetism (AM) in collinear magnetic systems is therefore highly desirable, yet remains a fundamental challenge. Floquet engineering via periodic driving provides a versatile paradigm for dynamically reshaping electronic landscapes and inducing exotic nonequilibrium quantum phases [36, 37, 38, 39, 40, 41, 42, 43, 39, 44, 45, 46, 47, 48, 49, 50, 51, 52]. While periodic light can effectively modify magnetic symmetries, the dynamical manipulation of spin-space-group symmetries to engineer higher-order odd-parity AM remains a pivotal and formidable frontier.
@@
Building on the $f$-wave AM state, on-resonant CPL driving with photon energies spanning half to the full band gap reshapes the electronic landscape. Through the optical Stark effect, photon-dressed Floquet-Bloch bands are shifted toward the Fermi level [76, 77, 78, 79], lifting spin degeneracies and triggering spin-layer-resolved TPTs characterized by band gap closing and reopening. The associated inversion of orbital characters provides an efficient means to redistribute the OAM and thereby modulate the OHE [80]. The OAM current generated by the OHE exerts an orbital torque upon injection into a ferromagnet. As illustrated in Fig.1(c), the Floquet AM manifests spin-layer locking, which generates layer-polarized torques and consequently drives counter-propagating electronic OAM currents under an external electric field. When the system is in a topologically nontrivial phase, the resulting dissipationless edge currents exhibit layer-locked orbital torque characteristics.
@@
To quantify the intrinsic orbital response, we evaluate the OHC by employing the Fermi-sea term of the Kubo formula within linear response theory [85, 86] $\sigma_{xy}^{L_{z}}\!=\!-e\!\int_{BZ}\frac{d^{2}\textbf{{k}}}{(2\pi)^{2}}\sum_{\nu}\!f_{\nu}(\textbf{{k}})\Omega_{\nu,xy}^{O,L_{z}}\!(\textbf{{k}})$ where $\Omega_{\nu,xy}^{O,L_{z}}\!(\textbf{{k}})$ is the OBC, given by $\Omega_{\nu,xy}^{O,L_{z}}\!(\textbf{{k}})\!=\!2\hbar\mathrm{Im}\sum_{\mu\neq\nu}\!\frac{\left\langle u_{\textbf{{k}}}^{\nu}\right|\mathcal{J}^{L_{z}}_{o,x}\left|u_{\textbf{{k}}}^{\mu}\right\rangle\left\langle u_{\textbf{{k}}}^{\mu}\right|\hat{v}_{y}\left|u_{\textbf{{k}}}^{\nu}\right\rangle}{(\varepsilon_{\textbf{{k}}}^{\nu}-\varepsilon_{\textbf{{k}}}^{\mu})^{2}}$, with $u_{\textbf{{k}}}^{\nu}$ representing the periodic part of the Bloch state associated with the energy $\varepsilon_{\textbf{{k}}}^{\nu}$. $f_{\nu}(\textbf{{k}})$ is the equilibrium Fermi distribution function and the velocity operator is denoted as $\hat{v}_{y}\!=\!\frac{1}{\hbar}\frac{\partial H_{\textbf{{k}}}}{\partial\textbf{{k}}_{y}}$. The OAM current operator is expressed as $\mathcal{J}^{L_{z}}_{o,x}\!=\!\frac{\left\{\hat{v}_{x},\hat{L}_{z}\right\}}{2}$, and for the Berry curvature (BC) calculations, this operator is substituted by the velocity operator $\hat{v}_{x}$. The OBC represents the geometric response of the momentum-space texture of OAM $\left\langle\textbf{{L}}_{z}(\textbf{{k}})\right\rangle\!=\!\left\langle u_{\textbf{{k}}}\right|\!\textbf{{L}}_{z}(\textbf{{k}})\!\left|u_{\textbf{{k}}}\right\rangle$. In the equilibrium bilayer, the OHC exhibits a plateau of $3.94~e/2\pi$ within the band gap [Fig.2(c)], with the OBC showing positive peaks at the $K$ and $K^{\prime}$ valleys. Under the coexistence of inversion $\mathcal{P}$ and time-reversal $\mathcal{T}$ symmetries, the OAM vanishes at the $K$ and $K^{\prime}$ points [Figs.S10-11] [84]. Furthermore, the emergence of floating edge states, in-gap modes, and localized charge distributions evidences the second-order topological insulator nature of the equilibrium system [Fig.S1] [87, 88, 89, 90, 84].
@@
where the positive (negative) sign distinguishes the lower (upper) layer, and $E_{c/v}$ are the eigenvalues of $H_{0}(\textbf{{k}},J_{0}(Aa)t_{ij}^{mn})$. Crucially, these splitting signs invert at the $K^{\prime}$ valley ($\textbf{{k}}\!\!\rightarrow\!\!-\textbf{{k}}$) , resulting in an odd-parity layer splitting [Figs.S2-S3] [84]. Interplaying with the intrinsic spin-layer locking, this effect induces an odd parity spin-splitting texture satisfying $E_{s}(\textbf{{k}})\!=\!E_{-s}(-\textbf{{k}})$. The emerging $f$-wave odd-parity AM is governed by the symmetry constraint $E_{s}(\textbf{{k}})\!\!=\!\!E_{-s}(\mathcal{O}^{-1}\textbf{{k}})$, where $\mathcal{O}$ represents the spatial component of the $C_{2}||S_{6z}]^{2}$. Fig.2(a) shows the spin-resolved 3D band structure of $H_{\rm{eff}}(\textbf{{k}})$, which demonstrates this light-induced higher-order odd-parity AM. Importantly, since AM originates entirely from light-induced symmetry breaking, time-periodic driving provides a general route to realize engineer AM states.
@@
Figure 2: (a) Spin-resolved 3D band structures for the $f$-wave odd-parity AM state described by $H_{\rm{eff}}(\textbf{{k}})$. (b) Spin-layer-resolved band structure near the $K$ and $K^{\prime}$ valley under RCPL with $\hbar\omega=0.10$ eV at light intensity $eA/\hbar=0.20~\mathrm{\AA }^{-1}$.(c) OHC as a function of energy for the equilibrium system and the nonequilibrium system driven by RCPL with different photon energies $\hbar\omega$ at $eA/\hbar=0.20~\mathrm{\AA }^{-1}$. (d) The orbital-resolved band structures near the $K$ valley under different photon energies $\hbar\omega=0.1,0.4,0.5$ eV at $eA/\hbar=0.20~\mathrm{\AA }^{-1}$.
@@
Notably, the photon-mediated processes encapsulated in $H_{F}(\textit{{k}},\omega)$ induce Floquet-Bloch sidebands that replicate the equilibrium electronic structure. Increasing the photon energy $\hbar\omega$ shifts these sidebands toward the Fermi level [Fig.S4] [84], providing a spectral signature for experimental detection via time-resolved angle-resolved photoemission spectroscopy (Tr-ARPES) [93]. In the sub-resonant regime where the driving energy remains well below the band-touching threshold (e.g., $\hbar\omega=0.10$ eV), the OHC is markedly amplified to nearly triple its equilibrium value, forming a robust plateau of $12.00~e/2\pi$. This surge originates from the coherent dressing of states, which triples the positive OBC peaks at the $K$ and $K^{\prime}$ valleys without necessitating a band gap closing [Fig.S10] [84]. Meanwhile, a zero anomalous Hall conductivity (AHC) plateau appears near the Fermi level [Figs.3(a) and (d)], determined by the momentum-space integration of the BC. Crucially, with increasing light intensity $eA/\hbar$, the spin degeneracy is further lifted, resulting in a significantly enhanced odd-parity spin splitting at the $K$ and $K^{\prime}$ valleys [Fig.2(b)]. Our findings reveal that coherent Floquet-dressing induces a massive amplification of pure orbital Hall transport and a field-tunable enhancement of odd-parity spin splitting even in the sub-resonance regime, establishing altermagnets as a reliable platform for petahertz-scale orbitronics.
@@
Figure 3: BC (left panel) and OBC (right panel) of the TB model in the $K$ valley under the RCPL with photon energy $\hbar\omega$ (a) =0.10=0.10 eV, (b) =0.40=0.40 eV and (c) =0.50=0.50 eV. Total and layer-resolved AHC under the RCPL with photon energy $\hbar\omega$ (d) =0.10=0.10 eV, (e) =0.40=0.40 eV and (f) =0.50=0.50 eV. The light intensity is $eA/\hbar=0.20~\mathrm{\AA ^{-1}}$. (g) OHC of the TB model under RCPL with $\hbar\omega=$ $0.10$ eV and 0.50 eV as a function of light intensity $eA/\hbar$.
@@
As the photon energy advances through the half-gap resonance, the system enters a topological regime where the emergence and subsequent gapping of spin-degenerate type-I nodal rings trigger a nontrivial orbital inversion [Fig.2(d)]. This evolution initiates a sequence of intricate topological phase transitions (TPTs) that fundamentally reconfigure the electronic landscape [Figs.S4–S11] [84]. Specifically, at $\hbar\omega\geq 0.20$ eV, orbital character exchange between the $E_{v}^{1}$ and $E_{c}^{-1}$ bands coincides with the first reopening of the global gap. This doubly degenerate band inversion generates positive OBC peaks at the $K$ and $K^{\prime}$ valleys, encircled by negative contributions, whereas the BC exhibits positive valley-centered peaks surrounded by positive rings [Fig.3(c)]. The resulting TPT establishes a quantized AHC plateau with a Chern number C=8C=-8 and an OHC of $8.05~e/2\pi$ [Figs.2(c) and 3(f)], with each layer contributing an equal $4~e^{2}/h$ to the anomalous Hall transport [94]. Significantly, the light-induced $f$-wave AM determines a valley-contrasting band inversion as light intensity further increases. A secondary inversion involving spin-polarized states (e.g., spin-up in the lower layer at $K$ valleys) reverses the valley-centered positive OBC peaks while leaving the surrounding negative rings intact [Fig.3(b)]. This evolution modifies the OHC to $6.07~e/2\pi$ and reconfigures the topological invariant to C=6C=-6. The quantized AHC reemerges with symmetric layer contributions of $3~e^{2}/h$, hallmarked by a BC landscape where negative valley peaks are enclosed by positive spectral rings [Fig.3(e)]. Collectively, light-induced AM underpins the co-evolution of discrete topological invariants and the orbital Hall response.
@@
Figure 4: (a) Crystal structure of the VSi2N4 bilayer. The gray region denotes the primitive cell. (b) The spin-resolved 3D band structures under RCPL. (c) Distribution of BC corresponding to the C=8C=-8 QAH phase. (d) Phase diagram of band gap as functions of photon energy $\hbar\omega$ and light intensity $eA/\hbar$.
Guided by symmetry considerations, we reveal that the interplay between photon energy and light intensity drives the VSi2N4 bilayer through a rich sequence of TPTs [Figs.4(d) and S15-S19] [84]. The phase diagram, mapped by the band gap in the ($\hbar\omega$, $eA/\hbar$) parameter space, exhibits four distinct regimes: a trivial insulator (I, $C_{l}=0$, $C_{u}=0$, $C_{\mathrm{tot}}=0$), a globally gapless phase (II), and two QAH phases with high Chern numbers (III, $C_{l}=-4$, $C_{u}=-4$, $C_{\mathrm{tot}}=-8$; IV, $C_{l}=-3$, $C_{u}=-3$, $C_{\mathrm{tot}}=-6$) [102], where $C_{l(u)}$ is the layer-resolved Chern number for the lower (upper) layer. In the trivial region I, the OHC is amplified to $7.01~e/2\pi$, approximately triple its equilibrium value. Upon entering the topological regime III, the OHC drops to $3.15~e/2\pi$, and further diminishes to $1.1~e/2\pi$ in region IV as valley-contrasting band inversions reshape the OBC distribution. Remarkably, the large topological invariants in these phases are underpinned by trigonal warping at the $K$ and $K^{\prime}$ valleys [103, 104, 105, 106], which manifests as three distinct BC peaks in the valence bands [Fig.4(c)]. Notably, this warping-induced proliferation of BC hot spots provides a geometric mechanism to elevate the Chern numbers. While $\mathcal{P}$ symmetry ensures mirror-symmetric BC distributions at $K$ and $K^{\prime}$, the light-driven warping expands the Floquet QAH phase space, facilitating the engineering of high-Chern-number states. The above mechanism can be observed in both trilayer TB models and VSi2N4 trilayer, where CPL can induce QAH phases with Chern numbers $C=\pm 12$ and $C=\pm 9$, and the OHC can reach up to $10.5~e/2\pi$ in VSi2N4. Moreover, reversing the CPL handedness flips the chirality of QAH phase while leaving the overall trend and sign of the OHC unchanged [Figs.S12 and S21-22] [84].

### 26 Mar 2026 v2

@@ Abstract
Introduction-Altermagnets provide a unique and significant platform for spin-polarized phenomena [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], distinguished by their nonrelativistic spin splitting in k-space protected by crystalline symmetries [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. Typically, altermagnets exhibit even-parity spin splitting, giving rise to Fermi surfaces with d-, g-, and i-wave symmetries [28, 29]. In contrast, magnets hosting unconventional odd-parity spin splitting (e.g., p- and f-wave types) are particularly significant due to their intimate connection with unconventional superfluidity [30, 31, 32]. However, odd-parity spin splitting has been largely limited to noncollinear magnets [33, 34, 35, 36, 37], whose complex magnetic textures are often fragile against external perturbations, restricting their practical applicability. Realizing odd-parity altermagnetism (AM) in collinear magnetic systems is therefore highly desirable, yet remains a fundamental challenge. Floquet engineering via periodic driving provides a versatile paradigm for dynamically reshaping electronic landscapes and inducing exotic nonequilibrium quantum phases [38, 39, 40, 41, 42, 43, 44, 45, 41, 46, 47, 48, 49, 50, 51, 52, 53, 54]. While periodic light can effectively modify magnetic symmetries, the dynamical manipulation of spin-space-group symmetries to engineer higher-order odd-parity AM remains a pivotal and formidable frontier [55, 56, 57, 58, 59, 60, 61, 62, 63].
@@
Building on the $f$-wave AM state, on-resonant CPL driving with photon energies spanning half to the full band gap reshapes the electronic landscape. Through the optical Stark effect, photon-dressed Floquet-Bloch bands are shifted toward the Fermi level [87, 88, 89, 90], lifting spin degeneracies and triggering spin-layer-resolved topological phase transitions (TPTs) characterized by band gap closing and reopening. The associated inversion of orbital characters provides an efficient means to redistribute the OAM and thereby modulate the OHE [91]. The OAM current generated by the OHE exerts an orbital torque upon injection into a ferromagnet. As illustrated in Fig.1(c), the Floquet AM manifests spin-layer locking, which generates layer-polarized torques and consequently drives counter-propagating electronic OAM currents under an external electric field. When the system is in a topologically nontrivial phase, the resulting dissipationless edge currents exhibit layer-locked orbital torque characteristics.
@@
To quantify the intrinsic orbital response, we evaluate the orbital Hall conductivity (OHC) by employing the Fermi-sea term of the Kubo formula within linear response theory [96, 97] $\sigma_{xy}^{L_{z}}\!=\!-e\!\int_{BZ}\frac{d^{2}\textbf{{k}}}{(2\pi)^{2}}\sum_{\nu}\!f_{\nu}(\textbf{{k}})\Omega_{\nu,xy}^{O,L_{z}}\!(\textbf{{k}})$ where $\Omega_{\nu,xy}^{O,L_{z}}\!(\textbf{{k}})$ is the OBC, given by $\Omega_{\nu,xy}^{O,L_{z}}\!(\textbf{{k}})\!=\!2\hbar\mathrm{Im}\sum_{\mu\neq\nu}\!\frac{\left\langle u_{\textbf{{k}}}^{\nu}\right|\mathcal{J}^{L_{z}}_{o,x}\left|u_{\textbf{{k}}}^{\mu}\right\rangle\left\langle u_{\textbf{{k}}}^{\mu}\right|\hat{v}_{y}\left|u_{\textbf{{k}}}^{\nu}\right\rangle}{(\varepsilon_{\textbf{{k}}}^{\nu}-\varepsilon_{\textbf{{k}}}^{\mu})^{2}}$, with $u_{\textbf{{k}}}^{\nu}$ representing the periodic part of the Bloch state associated with the energy $\varepsilon_{\textbf{{k}}}^{\nu}$. $f_{\nu}(\textbf{{k}})$ is the equilibrium Fermi distribution function and the velocity operator is denoted as $\hat{v}_{y}\!=\!\frac{1}{\hbar}\frac{\partial H_{\textbf{{k}}}}{\partial\textbf{{k}}_{y}}$. The OAM current operator is expressed as $\mathcal{J}^{L_{z}}_{o,x}\!=\!\frac{\left\{\hat{v}_{x},\hat{L}_{z}\right\}}{2}$, and for the Berry curvature (BC) calculations, this operator is substituted by the velocity operator $\hat{v}_{x}$. The OBC represents the geometric response of the momentum-space texture of OAM $\left\langle\textbf{{L}}_{z}(\textbf{{k}})\right\rangle\!=\!\left\langle u_{\textbf{{k}}}\right|\!\textbf{{L}}_{z}(\textbf{{k}})\!\left|u_{\textbf{{k}}}\right\rangle$. In the equilibrium bilayer, the OHC exhibits a plateau of $3.94~e/2\pi$ within the band gap [Fig.2(c)], with the OBC showing positive peaks at the $K$ and $K^{\prime}$ valleys. Under the coexistence of inversion $\mathcal{P}$ and time-reversal $\mathcal{T}$ symmetries, the OAM vanishes at the $K$ and $K^{\prime}$ points [Figs.S10-11] [95]. Furthermore, the emergence of floating edge states, in-gap modes, and localized charge distributions evidences the second-order topological insulator nature of the equilibrium system [Fig.S1] [98, 99, 100, 101, 95].
@@
where the positive (negative) sign distinguishes the lower (upper) layer, and $E_{c/v}$ are the eigenvalues of $H_{0}(\textbf{{k}},J_{0}(Aa)t_{ij}^{mn})$. Crucially, these splitting signs invert at the $K^{\prime}$ valley ($\textbf{{k}}\!\!\rightarrow\!\!-\textbf{{k}}$) , resulting in an odd-parity layer splitting [Figs.S2-S3] [95]. Interplaying with the intrinsic spin-layer locking, this effect induces an odd parity spin-splitting texture satisfying $E_{s}(\textbf{{k}})\!=\!E_{-s}(-\textbf{{k}})$. The emerging $f$-wave odd-parity AM is governed by the symmetry constraint $E_{s}(\textbf{{k}})\!\!=\!\!E_{-s}(\mathcal{O}^{-1}\textbf{{k}})$, where $\mathcal{O}$ represents the spatial component of the $C_{2}||S_{6z}]^{2}$. Fig.2(a) shows the spin-resolved 3D band structure of $H_{\rm{eff}}(\textbf{{k}})$, which demonstrates this light-induced higher-order odd-parity AM. Importantly, since AM originates entirely from light-induced symmetry breaking, time-periodic driving provides a general route to realize and engineer AM states.
@@
Figure 2: (a) Spin-resolved 3D band structures for the $f$-wave odd-parity AM state described by $H_{\rm{eff}}(\textbf{{k}})$. (b) Spin-layer-resolved band structure near the $K$ and $K^{\prime}$ valley under CPL with $\hbar\omega=0.10$ eV at light intensity $eA/\hbar=0.20~\mathrm{\AA }^{-1}$.(c) OHC as a function of energy for the equilibrium system and the nonequilibrium system driven by CPL with different photon energies $\hbar\omega$ at $eA/\hbar=0.20~\mathrm{\AA }^{-1}$. (d) The orbital-resolved band structures near the $K$ valley under different photon energies $\hbar\omega=0.1,0.4,0.5$ eV at $eA/\hbar=0.20~\mathrm{\AA }^{-1}$.
@@
Notably, the photon-mediated processes encapsulated in $H_{F}(\textit{{k}},\omega)$ induce Floquet-Bloch sidebands that replicate the equilibrium electronic structure. Increasing the photon energy $\hbar\omega$ shifts these sidebands toward the Fermi level [Fig.S4] [95], providing a spectral signature for experimental detection via time-resolved angle-resolved photoemission spectroscopy (Tr-ARPES) [104]. In the sub-resonant regime where the driving energy remains well below the band-touching threshold (e.g., $\hbar\omega=0.10$ eV), the OHC is markedly amplified to nearly triple its equilibrium value, forming a robust plateau of $12.00~e/2\pi$. This surge originates from the coherent dressing of states, which triples the positive OBC peaks at the $K$ and $K^{\prime}$ valleys without necessitating a band gap closing [Fig.S10] [95]. Meanwhile, a zero anomalous Hall conductivity (AHC) plateau appears near the Fermi level [Figs.3(a) and (d)], determined by the momentum-space integration of the BC. Crucially, with increasing light intensity $eA/\hbar$, the spin degeneracy is further lifted, resulting in a significantly enhanced odd-parity spin splitting at the $K$ and $K^{\prime}$ valleys [Fig.2(b)]. The massive, field-tunable enhancement of orbital Hall transport and odd-parity spin splitting via coherent Floquet-dressing establishes altermagnets as a reliable foundation for petahertz orbitronics.
@@
Figure 3: BC (left panel) and OBC (right panel) of the TB model in the $K$ valley under the CPL with photon energy $\hbar\omega$ (a) =0.10=0.10 eV, (b) =0.40=0.40 eV and (c) =0.50=0.50 eV. Total and layer-resolved AHC under the CPL with photon energy $\hbar\omega$ (d) =0.10=0.10 eV, (e) =0.40=0.40 eV and (f) =0.50=0.50 eV. The light intensity is $eA/\hbar=0.20~\mathrm{\AA ^{-1}}$. (g) OHC of the TB model under CPL with $\hbar\omega=$ $0.10$ eV and 0.50 eV as a function of light intensity $eA/\hbar$.
@@
As the photon energy advances through the half-gap resonance, the system enters a topological regime where the emergence and subsequent gapping of spin-degenerate type-I nodal rings trigger a nontrivial orbital inversion [Fig.2(d)]. This evolution initiates a sequence of intricate TPTs that fundamentally reconfigure the electronic landscape [Figs.S4–S11] [95]. Specifically, at $\hbar\omega\geq 0.20$ eV, orbital character exchange between the $E_{v}^{1}$ and $E_{c}^{-1}$ bands coincides with the first reopening of the global gap. This doubly degenerate band inversion generates positive OBC peaks at the $K$ and $K^{\prime}$ valleys, encircled by negative contributions, whereas the BC exhibits positive valley-centered peaks surrounded by positive rings [Fig.3(c)]. The resulting TPT establishes a quantized AHC plateau with a Chern number C=8C=-8 and an OHC of $8.05~e/2\pi$ [Figs.2(c) and 3(f)], with each layer contributing an equal $4~e^{2}/h$ to the anomalous Hall transport [105]. Significantly, the light-induced $f$-wave AM determines a valley-contrasting band inversion as light intensity further increases. This secondary inversion involving spin-polarized states (e.g., spin-up in the lower layer at $K$ valleys) reverses the valley-centered positive OBC peaks while leaving the surrounding negative rings intact [Fig.3(b)]. This evolution modifies the OHC to $6.07~e/2\pi$ and reconfigures the topological invariant to C=6C=-6. The quantized AHC reemerges with symmetric layer contributions of $3~e^{2}/h$, hallmarked by a BC landscape where negative valley peaks are enclosed by positive spectral rings [Fig.3(e)]. Collectively, light-induced AM underpins the co-evolution of discrete topological invariants and the orbital Hall response.
@@
Figure 4: (a) Crystal structure of the VSi2N4 bilayer. The gray region denotes the primitive cell. (b) The spin-resolved 3D band structures under CPL. (c) Distribution of BC corresponding to the C=8C=-8 QAH phase. (d) Phase diagram of band gap as functions of photon energy $\hbar\omega$ and light intensity $eA/\hbar$.
Guided by symmetry considerations, we reveal that the interplay between photon energy and light intensity drives the VSi2N4 bilayer through a rich sequence of TPTs [Figs.4(d) and S15-S19] [95]. The phase diagram, mapped by the band gap in the ($\hbar\omega$, $eA/\hbar$) parameter space, exhibits four distinct regimes: a trivial insulator (I, $C_{l}=0$, $C_{u}=0$, $C_{\mathrm{tot}}=0$), a globally gapless phase (II), and two QAH phases with high Chern numbers (III, $C_{l}=-4$, $C_{u}=-4$, $C_{\mathrm{tot}}=-8$; IV, $C_{l}=-3$, $C_{u}=-3$, $C_{\mathrm{tot}}=-6$) [113], where $C_{l(u)}$ is the layer-resolved Chern number for the lower (upper) layer. In the trivial region I, the OHC is amplified to $7.01~e/2\pi$, approximately triple its equilibrium value. Upon entering the topological regime III, the OHC drops to $3.15~e/2\pi$, and further diminishes to $1.1~e/2\pi$ in region IV as valley-contrasting band inversions reshape the OBC distribution. Remarkably, the large topological invariants in these phases are underpinned by trigonal warping at the $K$ and $K^{\prime}$ valleys [114, 115, 116, 117], which manifests as three distinct BC peaks in the valence bands [Fig.4(c)]. Notably, this warping-induced proliferation of BC hot spots provides a geometric mechanism to elevate the Chern numbers. While $\mathcal{P}$ symmetry ensures mirror-symmetric BC distributions at $K$ and $K^{\prime}$, the light-driven warping expands the Floquet QAH phase space, facilitating the engineering of high-Chern-number states. The above mechanism can be observed in both trilayer TB models and VSi2N4 trilayer, where CPL can induce QAH phases with Chern numbers $C=\pm 12$ and $C=\pm 9$, and the OHC can reach up to $10.5~e/2\pi$ in VSi2N4 [Figs.S12 and S21-22] [95]. Moreover, reversing the CPL handedness flips the chirality of QAH phase while leaving the overall trend and sign of the OHC unchanged.